OpenStudy (anonymous):

What is the length of the segment that joins the points: (-2, 3) and (-8, -7)? Write your answer in simplified radical form

OpenStudy (mathstudent55):

\(\bf \large \color{red}{The ~Distance ~Formula}\) The distance, d, between points \((x_1, y_1)\) and \((x_2, y_2) \) is \(d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \) Choose one point to be \((x_1, y_1)\) and the other point to be \((x_2, y_2) \) and substitute the coordinates in the formula to find d.

OpenStudy (anonymous):

Is the answer square root of 136?

OpenStudy (mathstudent55):

Let's try it: \(d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\) \(d = \sqrt{(-8 - (-2))^2 + (-7 - 3)^2}\) \(d = \sqrt{(-6)^2 + (-10)^2}\) \(d = \sqrt{36 + 100}\) \(d = \sqrt{136}\) You are correct, it is \( \sqrt{136} \). Now we need to see if we can simplify the radical.

OpenStudy (mathstudent55):

Does 136 have a perfect-square factor?

OpenStudy (anonymous):

34 square root 2?

OpenStudy (mathstudent55):

\(\sqrt{136} = \sqrt{4 \times 34} = 2\sqrt{34}\)

OpenStudy (mathstudent55):

Sorry, please notice the correct answer. The 34 stays in the square root. The square root of 4 comes out as 2.

OpenStudy (anonymous):

Thank you! and while youre here can you tell me if I got this other question right? the question was: What is the midpoint of the segment that joins the endpoints (-3, 12) and (-7, -20)? I got (-5, -4)

OpenStudy (mathstudent55):

Correct. There all you need to do is average the two x-coordinates and average the two y-coordinates which you did correctly. x-coordinate: (-3 + (-7))/2 = -10/2 = -5 y-coordinate: (12 + (-20))/2 = -8/2 = -4 Answer: (-5, -4), exactly what you got. Great job!

OpenStudy (anonymous):

Thank you! I have two other questions I am unsure of that I will post. Thanks again!!